Question:
The area of a rectangle is 4m^2 and its length is 3m longer that its width. Find the dimensions of the rectangle.

Answer by troyapplehelen(46)

A=w.l
4=x.(x+3)
0=x^2+3x-4
=(x-1)(x+4) x= 1 or -4, h/w, there is no such - number of length. so,
x, which is width, is 1 and its length is 4(=1+3).
The Answer width is 1 m, length is 4 m.